OpenStudy (anonymous):

Integral calculus help please Prove the third law of logarithms. Hint: Start by showing that both sides of the equation have the same derivative.

OpenStudy (anonymous):

do you have an example or is it just asking you to prove it?

OpenStudy (ranga):

http://www.youtube.com/watch?v=LgeRUr8ALGI

OpenStudy (ranga):

you are welcome.

OpenStudy (rational):

I think you need to use mean value theorem

OpenStudy (rational):

say \(f(x) = \ln (x^r)\) and \(g(x) = r \ln (x)\) \(f'(x) = \dfrac{1}{x^r} rx^{r-1} = \dfrac{r}{x}\) \(g'(x) = \dfrac{r}{x}\) That means \(f'(x) = g'(x)\)

OpenStudy (rational):

thats the first step only ^

OpenStudy (rational):

Any idea how to conclude ?

OpenStudy (ranga):

Hint: Start by showing that both sides of the equation have the same derivative. Didn't read the full question! @rational is correct.